OpenStudy (elise_a18):

Choose the polynomial that is written in standard form. 2x3y2 + 4x4y + 8x7 -3x4y2 + 4x4y5 + 10x2 6xy2 + 4x3y8 + 9x2 -2x5y3 + 4x2y2 + x2

OpenStudy (elise_a18):

the numbers after the variables are exponents ex.) 23x5 (5=exp.)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

go from highest degree to lowest degree is the sum of the exponents so for example the degree of \(-3x^4y^2=4+2=6\)

OpenStudy (elise_a18):

So it's -3xy^6?

OpenStudy (misty1212):

hmm no

OpenStudy (misty1212):

\[ -2x^5y^3 + 4x^2y^2 + x^2\] degrees of the terms are 8,4,2 from highest to lowest

OpenStudy (elise_a18):

So the one with terms going highest to lowest is the answer??? O-o

OpenStudy (misty1212):

yes

OpenStudy (elise_a18):

So D then?

OpenStudy (misty1212):

yes

OpenStudy (anonymous):

if anything i know a good sight that helps u is https://www.mathway.com/ its really good

OpenStudy (elise_a18):

Thanks so much! I love how you helped me get the answer without dragging it out. I like to learn how to do problems but a lot of the time the explanation gets too long or someone just hands me the answer. Thanks again! :)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

i suck at math lol v.v

OpenStudy (elise_a18):

Thanks for the link Gangsta XP

OpenStudy (anonymous):

np lol